please help me solve this equation !!
0.7(3x+6)=1.1-(x+2)

Answer by stanbon(57307)

0.7(3x+6)=1.1-(x+2)
---
Multiply both side by 10 to get:
7(3x+6) = 11 - 10(x+2)
---
21x+42 = 11 -10x-20
31x = -9-42
31x = -51
x = -51/31
=============
Cheers,
Stan H.
=============